OpenStudy (anonymous):

what is the radical of 338 simplified?

OpenStudy (mathstudent55):

Can you find the prime factors of 338?

OpenStudy (anonymous):

no ...I don't remember prime factorization..

OpenStudy (mathstudent55):

To do prime factorization, take the number, and divide it by the prime numbers starting with 2. Divide as many times by each prime number as possible before you move to the next prime number. The first several prime numbers are: 2, 3, 5, 7,11, 13, 17, 19, 23, 29

OpenStudy (mathstudent55):

338 is even. What is 338/2 = ?

OpenStudy (anonymous):

169

OpenStudy (mathstudent55):

Great. 169 is odd, so it's not divisible by 2. Is it divisible by 3?

OpenStudy (anonymous):

2*169=338

OpenStudy (mathstudent55):

We had that. Now take 169. It's not divisible by 2 bec its odd. We need to see which is the next prime number that 169 is divisible by. Is 169 divisible by 3?

OpenStudy (anonymous):

no..

OpenStudy (mathstudent55):

The quickest way to find out is to use a calculator. If 169/3 = a whole number, then it is divisible. If it is not a whole number, then it's not divisible.

OpenStudy (mathstudent55):

Good. The next prime is 5. 169 does not end in 0 or 5, so it's not divisible by 5. Try 7 next.

OpenStudy (anonymous):

no it doesn't go into 7

OpenStudy (mathstudent55):

Ok, move on to 11. If not, then try 13.

OpenStudy (anonymous):

13 works

OpenStudy (anonymous):

thank you!